# CCDM: Continuous Conditional Diffusion Models for Image Generation

### v2

@@ I Introduction
Continuous Conditional Generative Modeling (CCGM), as depicted in Fig. 1, aims to model the probability distribution of images conditioned on scalar continuous variables. These variables, commonly referred to as regression labels, can encompass a diverse range of values, including angles, ages, counts, temperatures, and more. However, CCGM faces significant challenges, particularly due to the scarcity of training samples for specific regression labels and the lack of an effective label input mechanism. These limitations adversely impact the performance of conventional Conditional Generative Adversarial Networks (cGANs) and Conditional Diffusion Models (CDMs), as demonstrated in [ding2023ccgan, ding2024turning].
@@
Diffusion models now lead image generation, consistently outperforming GANs in quality [dhariwal2021diffusion, ho2021classifier, rombach2022high, peebles2023scalable], making their adaptation to CCGM a promising path to surpass CcGANs [ding2023ccgan, ding2024turning]. However, existing Conditional Diffusion Models (CDMs) are designed for discrete conditions like class labels and text, making them unsuitable for regression labels in CCGM due to numerous limitations: their diffusion processes lack an explicit condition incorporation mechanism, leading to suboptimal performance, particularly for label inconsistency and poor visual fidelity; their label input mechanisms require modification for regression labels; the commonly used noise prediction loss and hyperparameter selection schemes are largely ineffective in CCGM; and the soft vicinal loss optimal for CcGANs does not translate well to CDMs. Consequently, integrating CDMs with CcGAN techniques faces significant challenges, underscoring an urgent need for developing novel diffusion processes, loss functions, label input mechanisms, and training/sampling strategies specially tailored to CCGM.
@@
  • In Section III-D, we present a multi-step algorithm for efficient conditional sampling, with key setups different from class-conditional or text-to-image scenarios. In Section III-E, we introduce DMD2-M, a one-step sampler inspired by DMD2 [yin2024improved], enhanced with advanced GAN architectures and training techniques to adapt it for the CCGM task. DMD2-M significantly accelerates sampling, with a slight trade-off in image quality.
@@
It is worth mentioning that the proposed CCDM substantially differs from a recent work CcDPM [zhao2024ccdpm] mainly in three aspects: (1) CcDPM focuses on non-image data, whereas high-dimensional image generation is significantly more challenging than non-image ones. Also, CcDPM was designed for multi-dimensional labels, lacking an efficient mechanism for scalar regression labels, whereas our CCDMs are specifically developed for image data with scalar continuous labels, making CCDM the first CDM tailored for the CCGM task. (2) CcDPM employs a noise prediction loss weighted by soft vicinity, while CCDM proposes a novel image denoising loss weighted by hard vicinity, which our ablation study in Section LABEL:sec:ablation demonstrates significantly better performance in CCGM tasks. (3) CcDPM’s diffusion processes are not directly tied to regression labels, whereas CCDM explicitly integrates regression labels into both forward and reverse processes, enhancing image quality. (4) CCDM provides a mathematically rigorous derivation of its modified diffusion processes and training loss in Appendix LABEL:sec:supp_detailed_derivations, which is absent in CcDPM. These distinctions underscore CCDM’s unique effectiveness in image generation with scalar continuous labels.
@@ II-B Diffusion Models
However, DDIM still typically requires hundreds of steps for sample generation, which remains time-consuming. To enable few-step or even one-step sampling, distillation-based methods are employed, with Improved Distribution Matching Distillation (DMD2) [yin2024improved] representing a state-of-the-art approach. DMD2 distills a trained diffusion model into a U-Net-based generator and reuses the U-Net for a discriminator, along with real/fake score functions, to compute distillation losses. The generator is then utilized for few-step or one-step sampling.
@@ III Methodology
In this section, we introduce CCDM, a novel approach specially designed to estimate the conditional distribution $p(\bm{x}^{0}|y)$ based on $N$ image-label pairs $\{(\bm{x}^{0}_{i},y_{i})\}_{i=1}^{N}$. CCDM introduces a new $y$-dependent diffusion processes, proposes a novel hard vicinal image denoising loss for training, adapts the CcGAN’s label input mechanism to diffusion models, and proposes two efficient sampling algorithms. The overall workflow is illustrated in Fig. 2.
Figure 2: The overall workflow of CCDM. Our method introduces $y$-dependent conditional diffusion processes and a novel hard vicinal image denoising loss (HVIDL) for U-Net training. To integrate the regression label $y$ into CCDM, we use two types of embeddings: a short embedding $\bm{h}_{y}^{s}$, encoded by CcGAN’s embedding network for U-Net injection, and a long embedding $\bm{h}_{y}^{l}$, encoded by the proposed covariance embedding network for forward diffusion and HVIDL. After training, we propose a multi-step sampling algorithm combining CFG [ho2021classifier] and DDIM [song2021denoising] for optimal image quality. Additionally, for fast sampling with a slight quality trade-off, we adapt DMD2 [yin2024improved] for CCGM and introduce DMD2-M for one-step sampling.
@@ III-B Covariance Embedding Network
Although the denoising U-Net for DDPM [ronneberger2015u, ho2020denoising] supports the sinusoidal positional embedding [ho2020denoising] and Gaussian Fourier projection [song2021scorebased] for encoding time steps $t$, Table LABEL:tab:ab_effect_embedding shows that these methods are ineffective for embedding regression labels $y$. As illustrated in Fig. 2, $y$ comprises two types of embeddings: a short embedding $\bm{h}_{y}^{s}$, which serves as input to the denoising U-Net, and a long embedding $\bm{h}_{y}^{l}$, which defines the modified diffusion processes and contributes to CCDM’s training loss and sampling algorithms. We utilize CcGAN’s embedding network $\phi(y)$ [ding2023ccgan] to convert the scalar $y$ into a 128-dimensional vector $\bm{h}^{s}_{y}$. However, $\phi(y)$ is unsuitable for computing $\bm{h}_{y}^{l}$ due to the significantly higher dimension of $\bm{h}_{y}^{l}$ compared to $\bm{h}_{y}^{s}$. To address this, we adapt CcGAN’s label embedding network to create the covariance embedding network, denoted as $\phi(y)^{\prime}$, to convert $y$ into $\bm{h}_{y}^{l}$. We employ a training scheme nearly identical to [ding2023ccgan] to train $\phi(y)^{\prime}$, with modifications to the architecture and a reduction in the training epochs of an auxiliary ResNet34 network to suit our requirements. For a more detailed explanation, please refer to Appendix LABEL:sec:supp_cov_label_embedding.
@@ III-C Hard Vicinal Image Denoising Loss
where $C$ is a constant and
@@
Eq.(19) represents an image denoising (aka $\bm{x}^{0}$-prediction) loss that is weighted by both the hard vicinal weight $W_{h}$ and the inverse of the $y$-dependent covariance matrix $\bm{H}_{y_{j}+\delta}^{-1}$. We minimize this loss for model fitting, with the detailed training procedure outlined in Algorithm LABEL:alg:training (Appendix LABEL:sec:supp_algorithms).
@@ III-D CFG and DDIM-based Multi-Step Sampling
where the initial $\bm{x}^{T^{\prime}}$ is randomly sampled from $\mathcal{N}(\bm{0},\bm{H}_{y})$ instead of a standard Gaussian $\mathcal{N}(\bm{0},\bm{I})$, making the sampling algorithm (Algorithm LABEL:alg:multistep_sampling in the Appendix) different from [song2021denoising].
@@ III-E Distillation-based One-Step Sampling
Even with DDIM, conditional sampling with CFG remains time-consuming, requiring hours or days to generate sufficient samples for evaluation. To improve efficiency, we attempted to distill the trained CCDM into a one-step generator using DMD2 [yin2024improved], but its design and training scheme proved ineffective for CCGM, causing severe mode collapse on RC-49 ($64\times 64$) (see Fig.LABEL:fig:vanilla_dmd2_failure_example in the Appendix). The U-Net generator in DMD2 was unnecessary for one-step sampling, and the shared feature extraction head between the discriminator and fake score function weakened discriminator performance. Additionally, DMD2’s reliance on vanilla GAN loss and outdated training techniques made it unsuitable for CCGM. To address these issues, we propose DMD2-M (shown in Fig.2), a modified version with several key improvements:
  • Network architecture: DMD2-M adopts SNGAN [miyato2018spectral] for both generator $G$ and discriminator $D$, using a 128-dimensional Gaussian noise vector for the generator instead of a high-dimensional noisy image. The discriminator is trained independently, without sharing a feature extraction head with the fake score function model $\hat{\bm{x}}_{\bm{\varphi}}^{0}$.
  • Loss functions: DMD2-M replaces DMD2’s vanilla GAN loss with the hinge GAN loss [miyato2018spectral] and integrates hard vicinal weights (Eq. (16)) into all training losses, helping to mitigate mode collapse in limited-data scenarios.
  • Training techniques: DMD2-M incorporates DiffAugment [zhao2020differentiable] into the distillation process to stabilize adversarial training and enhance performance.

### v3

@@ I Introduction
Continuous Conditional Generative Modeling (CCGM), as depicted in Fig. 1, aims to model the probability distribution of images conditioned on scalar continuous variables. These variables, commonly referred to as regression labels, can encompass a diverse range of values, including angles, ages, counts, temperatures, coordinates, and more. However, CCGM faces significant challenges, particularly due to the scarcity of training samples for specific regression labels and the lack of an effective label input mechanism. These limitations adversely impact the performance of conventional Conditional Generative Adversarial Networks (cGANs) and Conditional Diffusion Models (CDMs), as demonstrated in [ding2023ccgan, ding2024turning].
@@
Diffusion models now lead image generation, consistently outperforming GANs in quality [dhariwal2021diffusion, ho2021classifier, rombach2022high, peebles2023scalable, esser2024scaling, labs2025flux1kontextflowmatching], making their adaptation to CCGM a promising path to surpass CcGANs [ding2023ccgan, ding2024turning]. However, existing Conditional Diffusion Models (CDMs) are designed for discrete conditions like class labels [dhariwal2021diffusion, ho2021classifier] or text descriptions [rombach2022high, peebles2023scalable, esser2024scaling, li2024g, qu2024discriminative, labs2025flux1kontextflowmatching], making them unsuitable for regression labels in CCGM due to numerous limitations: their diffusion processes lack an explicit condition incorporation mechanism, leading to suboptimal performance, particularly for label inconsistency and poor visual fidelity; their label input mechanisms require modification for regression labels; the commonly used noise prediction loss and hyperparameter selection schemes are largely ineffective in CCGM; and the soft vicinal loss optimal for CcGANs does not translate well to CDMs. Consequently, integrating CDMs with CcGAN techniques faces significant challenges, underscoring an urgent need for developing novel diffusion processes, loss functions, label input mechanisms, and training/sampling strategies specially tailored to CCGM.
@@
  • In Section III-D, we employ a multi-step sampling algorithm optimized for CCGM, with configurations distinct from class-conditional or text-to-image generation. In Section III-E, we introduce DMD2-M, a one-step sampler inspired by DMD2 [yin2024improved], enhanced with advanced GAN architectures and training techniques to adapt it for the CCGM task. DMD2-M significantly accelerates sampling, with a slight trade-off in image quality.
@@
It is worth mentioning that the proposed CCDM substantially differs from a recent work CcDPM [zhao2024ccdpm] in four aspects:
  • CcDPM focuses on non-image data, whereas high-dimensional image generation is significantly more challenging than non-image ones. Also, CcDPM was designed for multi-dimensional labels, lacking an efficient mechanism for scalar regression labels, whereas our CCDMs are specifically developed for image data with scalar continuous labels, making CCDM the first CDM tailored for the CCGM task.
  • CcDPM employs a noise prediction loss weighted by soft vicinity, while CCDM proposes a novel image denoising loss weighted by hard vicinity, which our ablation study in Section LABEL:sec:ablation demonstrates significantly better performance in CCGM tasks.
  • CcDPM’s diffusion processes are not directly tied to regression labels, whereas CCDM explicitly integrates regression labels into both forward and reverse processes, enhancing image quality.
  • CCDM provides a mathematically rigorous derivation of its modified diffusion processes and training loss in Appendix LABEL:sec:supp_detailed_derivations, which is absent in CcDPM. These distinctions underscore CCDM’s unique effectiveness in image generation with scalar continuous labels.
@@ II-B Diffusion Models
However, DDIM still typically requires hundreds of steps for sample generation, which remains time-consuming. To enable few-step or even one-step sampling, distillation-based methods are employed, with Improved Distribution Matching Distillation (DMD2) [yin2024improved] representing a state-of-the-art approach. DMD2 comprises a generator, a real score function, a fake score function, and a discriminator, all modeled with convolutional neural networks. The real score function is a fixed pre-trained U-Net from the diffusion model, while the generator and fake score function, also based on this U-Net, are trainable. The discriminator’s feature extraction network shares weights with the encoder of the fake score function and is followed by fully connected layers that output the probability of an input image being real. The training process for DMD2 alternates between two steps: first, optimizing the generator using a distribution matching loss combined with the vanilla GAN loss [goodfellow2014generative]; second, training the fake score function and the discriminator with an image denoising loss and the vanilla GAN loss. Once training is complete, the generator is used for image generation. Notably, DMD2 multiplies the GAN losses for both the discriminator and generator by two positive weights, $w_{D}$ and $w_{G}$, which are treated as hyperparameters.
@@
II-C Text-to-Image Generation
Text-to-image (T2I) generation similarly involves conditional distribution estimation, driving advances in modern diffusion models [rombach2022high, peebles2023scalable, esser2024scaling, labs2025flux1kontextflowmatching]. However, as noted in [ding2023ccgan, ding2024turning] and demonstrated in our experimental results (Section IV), current T2I architectures like Stable Diffusion [rombach2022high] exhibit critical limitations in CCGM tasks. First, despite their ability to produce realistic images from prompts, these models inherently lack the architectural capacity to address the challenge of scarce or even zero training data for specific regression labels. Second, their dependence on large text encoders like CLIP [radford2021learning] for prompt embedding—significantly more complex than CcGAN’s compact 5-layer perceptron $\phi(y)$—introduces substantial and unnecessary memory overhead in CCGM applications.
Recent advances in text-to-image alignment [qu2024discriminative, li2024g] have explored quantitative prompt conditioning, though primarily focusing on discrete counting tasks. These approaches, while making progress in their specific domains, prove inadequate for addressing CCGM requirements. A key limitation lies in their inability to properly model the conditional distribution $p(\bm{x}|y)$ for regression labels or text containing such labels. The current frameworks are typically constrained to processing discrete numerical values (e.g., “Two girls sitting on the grass”) rather than continuous regression labels such as angular measurements. More significantly, these methods inherit the same fundamental limitations as general T2I diffusion models (including Stable Diffusion) when applied to CCGM tasks, failing to overcome the core challenges of conditional generation in this domain.
Consequently, quantitative prompt-conditioned image generation remains a largely unexplored area of research.
@@ III Methodology
In this section, we introduce CCDM, a novel approach specially designed to estimate the conditional distribution $p(\bm{x}^{0}|y)$ based on $N$ image-label pairs $\{(\bm{x}^{0}_{i},y_{i})\}_{i=1}^{N}$. CCDM introduces a new $y$-dependent diffusion processes, proposes a novel hard vicinal image denoising loss for training, adapts the CcGAN’s label input mechanism to diffusion models, and proposes two efficient sampling algorithms. The overall workflow is depicted in Fig. 2, where novel or customized components are highlighted in red.
Figure 2: The overall workflow of CCDM. Our method introduces $y$-dependent conditional diffusion processes and a novel hard vicinal image denoising loss (HVIDL) for U-Net training. To integrate the regression label $y$ into CCDM, we use two types of embeddings: a short embedding $\bm{h}_{y}^{s}$, encoded by CcGAN’s embedding network for U-Net injection, and a long embedding $\bm{h}_{y}^{l}$, encoded by the proposed covariance embedding network for forward diffusion and HVIDL. For inference, we employ a multi-step sampling algorithm combining CFG [ho2021classifier] and DDIM [song2021denoising], specially optimized for CCGM. We further introduce DMD2-M, an adaptation of DMD2 [yin2024improved] for one-step CCGM sampling with minimal quality trade-off.
@@ III-B Covariance Embedding Network
Although the denoising U-Net for DDPM [ronneberger2015u, ho2020denoising] supports the sinusoidal positional embedding [ho2020denoising] and Gaussian Fourier projection [song2021scorebased] for encoding time steps $t$, Table LABEL:tab:ab_effect_embedding shows that these methods are ineffective for embedding regression labels $y$. As illustrated in Fig. 2, $y$ comprises two types of embeddings: (1) a short embedding $\bm{h}_{y}^{s}\in\mathbb{R}^{128}$, which serves as input to the denoising U-Net, and (2) a long embedding $\bm{h}_{y}^{l}\in\mathbb{R}^{(C\times W\times H)}$, which modulates the diffusion processes and contributes to both CCDM’s training loss and sampling algorithms, where $C$, $W$, and $H$ denote the number of channels, width, and height of the input images, respectively. We utilize CcGAN’s embedding network $\phi(y)$ [ding2023ccgan] to convert the scalar $y$ into a 128-dimensional vector $\bm{h}^{s}_{y}$. However, $\phi(y)$ is unsuitable for computing $\bm{h}_{y}^{l}$ due to the significantly higher dimension of $\bm{h}_{y}^{l}$ compared to $\bm{h}_{y}^{s}$. To address this, we adapt CcGAN’s label embedding network to create the covariance embedding network, denoted as $\phi^{\prime}(y)$, to convert $y$ into $\bm{h}_{y}^{l}$. We employ a training scheme nearly identical to [ding2023ccgan] to train $\phi^{\prime}(y)$, with modifications to the architecture and a reduction in the training epochs of an auxiliary ResNet34 network to suit our requirements. For a more detailed explanation, please refer to Appendix LABEL:sec:supp_cov_label_embedding.
@@ III-C Hard Vicinal Image Denoising Loss
where $\sigma_{\delta}$ is the bandwidth for Kernel Density Estimation (KDE) of $p(y)$, $C$ is a constant, and
@@
Eq.(19) represents an image denoising (aka $\bm{x}^{0}$-prediction) loss that is weighted by both the hard vicinal weight $W_{h}$ and the inverse of the $y$-dependent covariance matrix $\bm{H}_{y_{j}+\delta}^{-1}$. We minimize this loss for model fitting, with the detailed training procedure outlined in Algorithm 1.
@@ III-D CFG and DDIM-based Multi-Step Sampling
where the initial $\bm{x}^{T^{\prime}}$ is randomly sampled from $\mathcal{N}(\bm{0},\bm{H}_{y})$ instead of a standard Gaussian $\mathcal{N}(\bm{0},\bm{I})$, making the sampling algorithm (Algorithm 2) different from [song2021denoising].
@@ III-E Distillation-based One-Step Sampling
Even with DDIM, conditional sampling with CFG remains time-consuming, often taking hours or days to generate sufficient evaluation samples. To enhance efficiency, we attempted to distill the trained CCDM into a one-step generator using DMD2 [yin2024improved], but its model design and training approach proved ineffective for our CCGM task. For instance, when applied to RC-49 ($64\times 64$), we observed severe mode collapse, as shown in Fig. 3 and Table LABEL:tab:ablation_dmd2m. Furthermore, one-step sampling does not require a U-Net architecture, and sharing the feature extraction head between the discriminator and fake score function harms the discriminator’s performance, as these models need distinct feature representations. Additionally, DMD2’s reliance on vanilla GAN loss and outdated training techniques makes it unsuitable for CCGM tasks, which often face data insufficiency and imbalance. To address these issues, we propose DMD2-M, an enhanced version of DMD2 with the following improvements:
  • Network architecture: DMD2-M adopts the SNGAN [miyato2018spectral] architecture for both generator $G$ and discriminator $D$, using a 128-dimensional Gaussian noise for the generator instead of a high-dimensional noisy image. The discriminator is trained independently and does not share a feature extraction head with the fake score function model.
  • Loss function: DMD2-M significantly improves upon DMD2 by adopting the hinge GAN loss [miyato2018spectral] in place of the vanilla GAN loss. Moreover, it incorporates the hard vicinal weight (Eq. (16)) into all training losses to help mitigate mode collapse in data-limited scenarios. Additionally, our proposed $y$-dependent covariance matrix $\bm{H}_{y}$ is integrated into the forward diffusion process for the computation of training losses.
  • Training technique: DMD2-M integrates DiffAugment [zhao2020differentiable] into the distillation process—a highly effective and user-friendly technique that stabilizes adversarial training and improves performance, especially in situations with limited training data.
@@
Figure 3: DMD2 [yin2024improved] suffers from severe mode collapse on RC-49 ($64\times 64$), demonstrating its inapplicability to the CCGM task.
Data: $N$ real image-label pairs $\Omega=\{(\bm{x}^{0}_{i},y_{i})\}_{i=1}^{N}$, $\tilde{N}$ distinct labels $\Upsilon=\{y_{[i]}\}_{i=1}^{\tilde{N}}$ arranged in ascending order from the training dataset;
Input: The pre-trained CcGAN’s embedding network $\phi(y)$ [ding2023ccgan]; The pre-trained covariance embedding network $\phi^{\prime}(y)$; Preset hard vicinity radius $\kappa$, KDE bandwidth $\sigma_{\delta}$, time steps $T$, condition drop probability $p_{\text{drop}}$, number of iterations $K$, and batch size $m$;
Result: A trained U-Net $\hat{\bm{x}}_{\bm{\theta}}^{0}$
1 for $k=1$ to $K$ do
2       Draw $m$ regression labels, denoted by $Y$, with replacement from $\Upsilon$;
3       Create a batch of target labels $Y^{\delta}=\{y_{i}+\delta|y_{i}\in Y,\delta\sim\mathcal{N}(0,\sigma_{\delta}^{2}),i=1,\dots,m\}$ ($\hat{\bm{x}}_{\bm{\theta}}^{0}$ is conditional on these labels) ;
4       Initialize $\Omega^{y_{i}+\delta}=\emptyset$ ;
5       for $i=1$ to $m$ do
6             Randomly choose an image-label pair $(\bm{x}^{0},y)\in\Omega$ satisfying $|y-y_{i}-\delta|\leq\kappa$ where $y_{i}+\delta\in Y^{\delta}$;
7             Draw $u\sim U(0,1)$;
8             if $u\geq p_{\text{drop}}$ then
9                   Update $\Omega^{y_{i}+\delta}=\Omega^{y_{i}+\delta}\cup(\bm{x}^{0},y_{i}+\delta)$;
10
11             else
12                   Update $\Omega^{y_{i}+\delta}=\Omega^{y_{i}+\delta}\cup(\bm{x}^{0},\emptyset)$;
13
14             end if
15
16       end for
17      Convert regression labels in $\Omega^{y_{i}+\delta}$ into their multi-dimensional representations $\bm{h}_{y_{i}+\delta}^{s}$ and $\bm{h}_{y_{i}+\delta}^{l}$ by applying the functions $\phi(y)$ and $\phi^{\prime}(y)$, respectively ;
18       Compute the $y$-dependent covariance matrix $\bm{H}_{y_{i}+\delta}$ via $\bm{H}_{y_{i}+\delta}=\text{diag}(\exp(-\bm{h}^{l}_{y_{i}+\delta}))$ ;
19       Update $\hat{\bm{x}}_{\bm{\theta}}^{0}$ using samples in the set $\Omega^{y_{i}+\delta}$ via gradient-based optimizers according to Eq. (19);
20
21 end for
Algorithm 1 Algorithm for Training CCDM
Input: The pre-trained U-Net $\hat{\bm{x}}_{\bm{\theta}}^{0}$; The pre-trained CcGAN’s embedding network $\phi(y)$ [ding2023ccgan]; The pre-trained covariance embedding network $\phi^{\prime}(y)$; The target regression label $y$; Preset hyperparameters, including the time steps $T^{\prime}$ for sampling, the conditional scale $\gamma$, and the variance schedule $\beta_{i}$’s;
Result: A fake image $\bm{x}^{0}_{g}$ with label $y$
1 Convert the regression label $y$ into its multi-dimensional representations $\bm{h}_{y}^{s}$ and $\bm{h}_{y}^{l}$ by applying the functions $\phi(y)$ and $\phi^{\prime}(y)$, respectively ;
2 Compute the $y$-dependent covariance matrix $\bm{H}_{y}$ via $\bm{H}_{y}=\text{diag}(\exp(-\bm{h}^{l}_{y}))$ ;
3 $t\leftarrow T^{\prime}$, $\bm{x}^{T^{\prime}}_{g}\sim\mathcal{N}(\bm{0},\bm{H}_{y})$ ;
4 while $t>0$ do
5       Compute $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$;
6       Compute $\tilde{\bm{x}}^{0}=(1-\gamma)\cdot\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t}_{g},t,\oslash)+\gamma\cdot\hat{\bm{x}}_{\bm{\theta}}^{0}(\bm{x}^{t}_{g},t,y)$;
7       Update $\bm{x}^{t-1}_{g}=\sqrt{\bar{\alpha}_{t-1}}\tilde{\bm{x}}^{0}+\sqrt{1-\bar{\alpha}_{t-1}}\frac{\bm{x}^{t}_{g}-\sqrt{\bar{\alpha}_{t}}\tilde{\bm{x}}^{0}}{\sqrt{1-\bar{\alpha}_{t}}}$;
8       Update $t\leftarrow t-1$;
9 end while
Algorithm 2 Multi-Step Sampling Algorithm for CCDM Utilizing CFG and DDIM